OpenStudy (anonymous):

solve using elimination: u+v+w=6, u+2v+2w=11, 2u+3v-4w=3.....i keep getting the wrong answers according to wolf ram and i am not sure why

OpenStudy (campbell_st):

label your equations (1), (2), (3) and reduce the 3 equations to 2 equations in 2 unknowns.. Eliminate u Equation (1) - Equation (2) gives -v -w = -5 (4) 2 x Equation (1) - Equations(3) gives -v + 6w = 9 (5) Eliminate v -1 x Equation (4) + equation (5) = 7w = 14 Substitute to get the other pronumerals. My solution is u =1, v = 3 and w = 2

OpenStudy (phi):

u+v+w=6, u+2v+2w=11, 2u+3v-4w=3 just the coefficients, in order u v w constant 1 1 1 6 1 2 2 11 2 3 -4 3 multiply the top row by -1 and add to the 2nd row ( we add -1 -1 -1 -6 to the 2nd row) 1 1 1 6 0 1 1 5 2 3 -4 3 multiply the top row by -2 and add to the 3rd row (we add -2 -2 -2 -12 to the 3rd row) 1 1 1 6 0 1 1 5 0 1 -6 -9 multiply the 2nd row by -1 and add to the 3rd row (add 0 -1 -1 -5 to the 3rd row) 1 1 1 6 0 1 1 5 0 0 -7 -14 backsolve (start at the bottom) w= -14/-7 = 2 v + 2= 5, v= 3 u+3+2=6, u = 1

OpenStudy (anonymous):

Answer : u = 1 v= 3 w = 2